Grouper Company owns a 7,700-acre tract of timber purchased in 2006 at a cost of $1,404 per acre. At the time of purchase, the land was estimated to have a value of $324 per acre without the timber. Grouper Company has not logged this tract since it was purchased. In 2020, Grouper had the timber cruised. The cruise (appraiser) estimated that each acre contained 8,640 board feet of timber. In 2020, Grouper built 10 miles of roads at a cost of $7,940 per mile. After the roads were completed, Grouper logged and sold 3,780 trees containing 920,000 board feet.

Collapse question part

(a)

Determine the cost of timber sold related to depletion for 2020. (Round intermediate calculations to 5 decimal places, e.g. 1.24654 and final answer to 0 decimal places, e.g. 5,125.)
Cost of timber sold $

b) If Grouper depreciates the logging roads on the basis of timber cut, determine the depreciation expense for 2020

Depreciation expense $_______

c) If Grouper plants five seedlings at a cost of $4 per seedling for each tree cut, how should Grouper treat the reforestation?

Grouper should capitalize the cost of $_________
Answer #1

Cost of timber per acre (without the land) = Total cost - Cost of land

                                                                         = 1404 - 324 = 1080

Cost of timber per feet = Cost of timber per acre/Feet of timber per acre

                                          = 1080/8640 = 0.125

a) Cost of timber sold = Cost of timber per feet X Feet sold

                                   = 0.125 X 920000 = $115,000

b) The useful life of the logging road is not given. Thus its depreciation for 2020 cannot be calculated.

c) Reforestation cost = No. of Trees cut X No. of seeds X Cost per seed

                                        = 3780 X 5 X 4 = $75,600

The benefit of these seeds will extend beyond the current financial period as the timber grown from these seeds will keep on increasing from year to year. Thus it is a capital expenditure. Grouper should capitalize the cost of $ 75,600